OpenStudy (anonymous):

x – 2y = 11 substitute above into formula and solve x = 1 + 3y

OpenStudy (anonymous):

x - 2y = 11 => x = 2y+11. plug that into second equation. 2y+11 = 1+ 3y => 11=1+y => y = 10. plug y into either of the equations. x=1+3y = 1+3(10) = 31. so x = 31,y=10

OpenStudy (anonymous):